OpenStudy (anonymous):

Solve the system by substitution. x + y = 5 y = x – 1

OpenStudy (solomonzelman):

Look at what y equals to in the second equation. See it? Substitute that instead of y into the first equation.

OpenStudy (solomonzelman):

x + y = 5 y = x – 1 x + (x-1) = 5 x=? SOLVE FOR X....

OpenStudy (anonymous):

i got three is that right

OpenStudy (solomonzelman):

Yes x=3! Can you solve for y?

OpenStudy (solomonzelman):

But, not you, @kimmybowlin I want @Ambbiiee to do it, since she is the one that asked.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is algebra right

OpenStudy (solomonzelman):

Yes, this is algebra 1.

OpenStudy (anonymous):

y-5=-2y+1 y+2y=1+5 3y=6 y=6/3 y=2

OpenStudy (anonymous):

ok i passed it a while ago and was trying to see if i still have my skills that were portrayed during that time. but thank you!!:)

OpenStudy (solomonzelman):

Yes, that's correct y=2... @Ambbiiee if you will, can you right the answer as an (x,y) coordinate? You absolutely don't have to do it, I just want to see you do it:)

OpenStudy (solomonzelman):

I am not sure what exactly are you doing here -:(

OpenStudy (anonymous):

how did you guys get 3

OpenStudy (anonymous):

x+(x-1)=5 2x-1=5 +1 +1 2x=6 /2 /2 x=3

OpenStudy (solomonzelman):

x + y = 5 y = x – 1 substitution x+(x-1)=5 x+x-1=5 2x-1=5 (add 1 to both sides)... 2x=6 x=3 x + y = 5 3 + y = 5 (subtract 3 from both sides) y = 2

OpenStudy (solomonzelman):

Makes sense @Ambbiiee , right?

OpenStudy (anonymous):

slowly coming to an understanding @SolomonZelman

OpenStudy (solomonzelman):

alright, I can explain it a little better, or give you more time....